OpenStudy (anonymous):

12 is what percent of 36

OpenStudy (anonymous):

\[\large 12=\frac{x}{100}\times 36\]

OpenStudy (anonymous):

\[\large \rightarrow 12=\frac{36}{100}x\]

OpenStudy (anonymous):

can you show me how to solve it

OpenStudy (anonymous):

\[\large \rightarrow 12 \cdot \frac{100}{36}=x\]

OpenStudy (anonymous):

x = 1200 ÷ 36

OpenStudy (anonymous):

thanks wanna help me with one more just to make sure i understand

OpenStudy (anonymous):

Ok.